# Large Language Model Situational Awareness Based Planning

### 26 Dec 2023 v1

@@ Abstract
This work pioneers evaluating emergent planning capabilities based on situational awareness in large language models. We contribute (i) novel benchmarks and metrics for standardized assessment; (ii) a unique dataset to spur progress; and (iii) demonstrations that prompting and multi-agent schemes significantly enhance planning performance in context-sensitive planning tasks. Positioning this within a situated agent and automated planning research, we highlight inherent reliability challenges–efficiently mapping world states to actions without environmental guidance remains open despite simulated domain advances. Although out-of-scope, limitations around validation methodology and data availability indicate exciting directions, including fine-tuning on expanded planning corpora and optimizations for triggering fast latent planning. By conclusively demonstrating current methods’ promise and limitations via rigorous comparison, we catalyze investigating reliable goal-directed reasoning for situated agents.
@@ 1 Introduction
This research shows LLMs can exhibit human-like planning capacities based on situational awareness. As Fig. 1 illustrates, when prompted for situational awareness based planning spanning perception, comprehension and projection [2] or when provided reasoning feedback, the LLM displays improved deductive reasoning requiring perspective-taking and contemplating potential outcomes. This study distinguishes itself from previous AI reasoning and planning research in significant ways. Much existing work evaluates task-specific planning or makes assumptions about predefined steps under strict conditions [7, 8, 9, 10, 11, 12]. Additionally, various agents reactively generate plans only after explicit instructions or demands [13, 14, 15, 16, 17, 18, 19, 20], rather than proactively. In contrast, our approach centres on assessing and enhancing models’ capacities for proactive situational planning when faced with real, open-world dilemmas. Without rigid constraints or environmental feedback, the models must leverage deductive logic to map conjectured actions while weighing consequences using only a descriptive starting scenario and prompts. By prioritizing latent planning that mirrors human cognition - namely situational awareness based rather than domain-limited planning - our method signifies a break from former suppositions and appraisals.
@@ 2.3 State-based Planning with feedback
Fig. 2 visually depicts the iterative process between $\text{LLM}_{gen}$ and $\text{LLM}_{eval}$. Consider a scenario where a young child attempts to touch a hot pot on an active stove, posing a safety risk. The housekeeper robot observing this scene starts planning appropriate interventions. First, the instruction-prompted $\text{LLM}_{gen}$ leverages its reasoning capacity to imagine potential outcomes and draft candidate FSM plans to mitigate harm. For instance, abruptly obstructing the child may startle the child, suggesting a gentle approach or showing toys instead. If the child refuses intervention and gets burned, emergency responses may become necessary. $\text{LLM}_{gen}$ passes its proposed plan $\hat{M}$ with the comparison plan - which is a human demo at the first round then the previous plan - to $\text{LLM}_{eval}$ for evaluative scoring and feedback $f$. $\text{LLM}_{gen}$ then incorporates this $f$ into the next round of planning. Over one or more proposal-evaluation iterations, the agents converge on a new optimal FSM plan $M^{*}$’s score surpasses the benchmark plan score that can morally and robustly handle edge cases through situational inference.
@@ 3.1 Evaluation Scenarios
We first collected a dataset of over 500 hazardous scenarios situated in home environments. From this dataset, we systematically sampled 24 representative vignettes showcasing common home hazards across four reasoning complexity levels, as categorized in Table 1 and Fig. 3. High-resolution textual descriptions of these 24 evaluation scenes were then produced using GPT-4[25]. Accompanying gold standard solutions formally encoding state-based hazard interventions were manually constructed and verified by experts for each scene. This standardized 24-scene benchmark with fluent human reasoning enables quantitative analysis of LLM planning capacities relative to human proficiency. The full dataset of over 500 hazardous home scenarios is provided to enable future fine-tuning and extension of models. By summarizing the key complexity dimensions and gold standard solutions for 24 systematically selected scenes, we aim to evaluate model planning capacities on a wide distribution of common situations requiring hazard intervention. The full dataset allows pushing beyond the scope of the defined benchmark towards continual learning of planning skills for household robots.
@@ 3.3 Evaluation dimensions
Seven scoring dimensions have been developed as shown in Table 3.1. These seven dimensions provide a comprehensive methodology for assessing the latent planning - FSM designs. Touching on coverage, complexity, safety, reusability, user experience and coherence, the framework allows evaluating structured completeness, validation requirements, real-world reliability, adaptability, human factors and solution integrity. Using these lenses together promotes designs that are robust, dependable, future-proof, ethical and aligned to specifications. The dimensions offer multi-faceted technical and operational insight. Scoring FSMs across seven key dimensions on a scale from 0 to 10 enables impartial quantitative evaluation of overall plan quality while revealing relative strengths and weaknesses to prioritize refinements. The overall score is calculated as the average mean of the seven dimension scores.
@@ 4.2 Impact of the SAP Prompt
An experiment is conducted to evaluate the effect of the SAP prompt on the quality of planning. As shown in Table 3, three LLMs - GPT-4, GPT-3.5, and Claude-2 - are tested with or without the SAP prompt on the benchmark scenarios across four complexity levels.
The RBS methodology is utilized, where models are compared pair-wise on each scenario and differentially ranked. Adding the SAP prompt leads to improved RBS scores for all three models compared to those not, indicating enhanced planning. Specifically, GPT-4 with the SAP prompt achieves the best overall RBS score of 1.21, significantly outperforming GPT-4 without prompts which has an RBS score of 2.04. Analysis of the reasoning level 3 scenarios - involving dangerous interactions with children, elderly, and pets - shows GPT-4 with the SAP prompt substantially exceeded the second-best model. This suggests the prompt provides particular value for complex, nuances planning situations requiring perspective-taking and contemplating potential outcomes (ablation studies see Appendix B).
The results demonstrate that incorporating prompts directing models to deeply consider relationships and iterative consequences significantly enhance latent planning capacities. By promoting coordination and foresight, deductions improve when reasoning about multi-agent safety hazards. However, in future works, tuning prompt design and training models directly on planning corpora can help strengthen reliable situational inference.
@@ 4.3 LLM Evaluators
The models are tasked with ranking FSM plans in pairs, groups of 4, and groups of 6. Rankings are then compared to expert human rank orders to quantify accuracy. Tests find both GPT-4 and Claude-2 could rank FSM pairs with 75.7% agreement to human ranking, evidencing reliability for comparative evaluation. However, accuracy drops significantly for ranking 4 or more FSMs. Table 5 and Fig. 4 show that GPT-4 and Claude-2 aligned best with human judgment when evaluating outputs from the models themselves. For example, GPT-4’s scoring of GPT-4 FSMs closely matches expert ranking. This demonstrates LLMs can provide sound comparative assessments, particularly for their own model family’s outputs. The experiments illustrate LLMs’ potential to serve as evaluators that mimic human appraisals of planning formalisms. By relying on relative rather than absolute assessments, variability is reduced. Ongoing work should explore training models expressly for evaluation to boost accuracy, as well as architecting dedicated scoring networks.
@@ 5 Conclusion
Our analysis provides initial evidence of improved planning from situational awareness prompts and multi-agent feedback. Experiments on a home hazard benchmark exhibit gains in reliability and humaneness when prompting relationship reasoning and consequence consideration, as exemplified by the planning outputs in Appendix C. Results specifically show models GPT-3.5, GPT-4, and Claude-2 demonstrate superior deduction and mapping with open-ended prompts relative to no prompts, especially for complex safety scenarios. Additionally, a complete closed-loop experiment proves that multi-agent collaboration boosts overall FSM quality. While advances are promising, efficiently condensing fluid cognition absent environmental feedback remains an open challenge. Important future work includes expanded planning corpus training and coupling deduction with grounded simulation. As capabilities progress, introduced techniques appear useful for instilling robustness. However, oversight guardrails remain vital for ensuring ethical alignment before real-world instantiation.
By assessing generative planning capacities and limitations, this research catalyzes investigating goal-directed reasoning at the intersection of deduction, awareness, and embodiment. It prudently steers progress in situated agents, balancing innovation with ethical caution.
Appendix A Appendix A
@@ A.2 Human Annotator Evaluation
Fig. 7: Models rankings distribution on 24 questions. Rankings are rounded.

### 01 Jan 2024 v2

@@ Abstract
This work pioneers evaluating emergent planning capabilities based on situational awareness in large language models. We contribute (i) novel benchmarks and metrics for standardised assessment; (ii) a unique dataset to spur progress; and (iii) demonstrations that are prompting and multi-agent schemes significantly enhance planning performance in context-sensitive planning tasks. Positioning this within a situated agent and automated planning research, we highlight inherent reliability challenges–efficiently mapping world states to actions without environmental guidance remains open despite simulated domain advances. Although out-of-scope, limitations around validation methodology and data availability indicate exciting directions, including fine-tuning on expanded planning corpora and optimizations for triggering fast latent planning. By conclusively demonstrating current methods’ promise and limitations via rigorous comparison, we catalyze investigating reliable goal-directed reasoning for situated agents.
@@ 1 Introduction
This research shows LLMs can exhibit human-like planning capacities based on situational awareness. As Fig. 1 illustrates, when prompted for situational awareness based planning spanning perception, comprehension and projection [2] or when provided reasoning feedback, the LLM displays improved deductive reasoning requiring perspective-taking and contemplating potential outcomes. This study distinguishes itself from previous AI reasoning and planning research in significant ways. Much existing work evaluates task-specific planning or makes assumptions about predefined steps under strict conditions [7, 8, 9, 10, 11, 12]. Additionally, various agents reactively generate plans only after explicit instructions or demands [13, 14, 15, 16, 17, 18, 19, 20], rather than proactively. In contrast, our approach centres on assessing and enhancing models’ capacities for proactive situational planning when faced with real, open-world dilemmas. Without rigid constraints or environmental feedback, the models must leverage deductive logic to map conjectured actions while weighing consequences using only a descriptive starting scenario and prompts. By prioritizing latent planning that mirrors human cognition, namely situational awareness based rather than domain-limited planning, our method signifies a break from former suppositions and appraisals.
@@ 2.3 State-based Planning with Feedback
Fig. 2 visually depicts the iterative process between $\text{LLM}_{gen}$ and $\text{LLM}_{eval}$. Consider a scenario where a young child attempts to touch a hot pot on an active stove, posing a safety risk. The housekeeper robot observing this scene starts planning appropriate interventions. First, the instruction-prompted $\text{LLM}_{gen}$ leverages its reasoning capacity to imagine potential outcomes and draft candidate FSM plans to mitigate harm. For instance, abruptly obstructing the child may startle the child, suggesting a gentle approach or showing toys instead. If the child refuses intervention and gets burned, emergency responses may become necessary. $\text{LLM}_{gen}$ passes its proposed plan $\hat{M}$ with the comparison plan, which is a human demo at the first round then is the previous plan, to $\text{LLM}_{eval}$ for evaluative scoring and feedback $f$. $\text{LLM}_{gen}$ then incorporates this $f$ into the next round of planning. Over one or more proposal-evaluation iterations, the agents converge on a new optimal FSM plan $M^{*}$’s score surpasses the benchmark plan score that can morally and robustly handle edge cases through situational inference.
@@ 3.1 Evaluation Scenarios
We first collected a dataset of over 500 hazardous scenarios situated in home environments. From this dataset, we systematically sampled 24 representative vignettes showcasing common home hazards across four reasoning complexity levels, as categorized in Table 1 and Fig. 3. High-resolution textual descriptions of these 24 evaluation scenes were then produced using GPT-4 [25]. Accompanying gold standard solutions formally encoding state-based hazard interventions were manually constructed and verified by experts for each scene. This standardized 24-scene benchmark with fluent human reasoning enables quantitative analysis of LLM planning capacities relative to human proficiency. The full dataset will be released. By summarizing the key complexity dimensions and gold standard solutions for 24 systematically selected scenes, we aim to evaluate model planning capacities on a wide distribution of common situations requiring hazard intervention. The full dataset allows pushing beyond the scope of the defined benchmark towards continual learning of planning skills for household robots.
@@ 3.3 Evaluation Dimensions
Seven scoring dimensions have been developed as shown in Table 3.1. These seven dimensions provide a comprehensive methodology for assessing the latent planning, FSM designs. Touching on coverage, complexity, safety, reusability, user experience and coherence, the framework allows evaluating structured completeness, validation requirements, real-world reliability, adaptability, human factors and solution integrity. Using these lenses together promotes designs that are robust, dependable, future-proof, ethical and aligned to specifications. The dimensions offer multi-faceted technical and operational insight. Scoring FSMs across seven key dimensions on a scale from 0 to 10 enables impartial quantitative evaluation of overall plan quality while revealing relative strengths and weaknesses to prioritize refinements. The overall score is calculated as the average mean of the seven dimension scores.
@@ 4.2 Impact of The SAP Prompt
An experiment is conducted to evaluate the effect of the SAP prompt on the quality of planning. As shown in Table 3, three LLMs, GPT-4, GPT-3.5, and Claude-2, are tested with or without the SAP prompt on the benchmark scenarios across four complexity levels.
The RBS methodology is utilized, where models are compared pair-wise on each scenario and differentially ranked. Adding the SAP prompt leads to improved RBS scores for all three models compared to those not, indicating enhanced planning. Specifically, GPT-4 with the SAP prompt achieves the best overall RBS score of 1.21, significantly outperforming GPT-4 without prompts which has an RBS score of 2.04. Analysis of the reasoning level 3 scenarios, involving dangerous interactions with children, elderly, and pets, shows GPT-4 with the SAP prompt substantially exceeded the second-best model. This suggests the prompt provides particular value for complex, nuances planning situations requiring perspective-taking and contemplating potential outcomes (ablation studies see Appendix B) [39, 40].
The results demonstrate that incorporating prompts directing models to deeply consider relationships and iterative consequences significantly enhance latent planning capacities. By promoting coordination and foresight, deductions improve when reasoning about multi-agent safety hazards.
@@ 4.3 LLM Evaluators
The models are tasked with ranking FSM plans in pairs, groups of 4, and groups of 6. Rankings are then compared to expert human rank orders to quantify accuracy. Tests find both GPT-4 and Claude-2 could rank FSM pairs with 75.7% agreement to human ranking, evidencing reliability for comparative evaluation. However, accuracy drops significantly for ranking 4 or more FSMs. Table 5 and Fig. 4 show that GPT-4 and Claude-2 aligned best with human judgment when evaluating outputs from the models themselves. For example, GPT-4’s scoring of GPT-4 FSMs closely matches expert ranking. This demonstrates LLMs can provide sound comparative assessments, particularly for their own model family’s outputs. The experiments illustrate LLMs’ potential to serve as evaluators that mimic human appraisals of planning formalisms. By relying on relative rather than absolute assessments, variability is reduced.
@@ 5 Conclusion
Our analysis provides initial evidence that situational awareness prompts and multi-agent feedback improve planning reliability and humaneness. Experiments on a home hazard benchmark show gains when prompting relationship reasoning and consequence consideration, as exemplified by the planning outputs in Appendix C. Results specifically demonstrate that GPT-3.5, GPT-4, and Claude-2 have superior deduction and mapping with open-ended prompts, especially for complex safety scenarios. Additionally, a complete closed-loop experiment proves that multi-agent collaboration boosts quality. While advances are promising, efficiently condensing fluid cognition absent environmental feedback remains an open challenge. Important future work includes expanded planning corpus training and coupling deduction with simulation. As capabilities progress, introduced techniques appear useful for instilling robustness. However, oversight guardrails remain vital for ensuring ethical alignment before real-world instantiation. By assessing generative planning capacities and limitations, this research steers progress at the intersection of deduction, awareness, and embodiment while balancing innovation with ethical caution.
Appendix A Appendix Supplement
@@ A.2 Human Annotator Evaluation
Fig. 7: Models rankings distribution on the 24 scenarios. Rankings are rounded.